OpenStudy (gazela1997):

PLEASE HELPP!!. FACTOR 3x^2+x-10

OpenStudy (shamim):

6x-5x=x

OpenStudy (shamim):

use it

OpenStudy (gazela1997):

?

OpenStudy (gazela1997):

@IMStuck

OpenStudy (imstuck):

Ok, let me look real quick. Ill be back in a sec.

OpenStudy (imstuck):

Your factors are ( x - 5/3) and (x + 2).

OpenStudy (gazela1997):

could you show me how?

OpenStudy (imstuck):

Yes. I had to use the quadratic formula. Do you know it?

OpenStudy (gazela1997):

YEP

OpenStudy (imstuck):

Ok, so if you know it, try it and let me know you got the same answers I did.

OpenStudy (gazela1997):

what would be x^2-6x?

OpenStudy (gazela1997):

i know the formul but idk how you used it

OpenStudy (imstuck):

This is the formula. What do you mean by what would be x^2 - 6x?

OpenStudy (gazela1997):

nvm about that

OpenStudy (imstuck):

\[\frac{ -b \pm \sqrt{b ^{2}-4(a)(c)} }{ 2a }\]We have values we can fill in there from our equation. In 3x^2 + x - 10, the a is 3, the b is 1, and the c is -10

OpenStudy (imstuck):

\[\frac{ -1\pm \sqrt{(1)^{2}-4(3)(-10)} }{ 2(3)}\]

OpenStudy (imstuck):

That gives us\[\frac{ -1\pm \sqrt{1+120} }{ 6 }\]or\[\frac{ -1\pm \sqrt{121} }{ 6 }\]or\[\frac{ -1\pm11 }{ 6 }\] -1 + 11/6=10/6=5/3. That's one solution. The other is -1-11/6=-12/6= -2. So your factors for that are (x + 2) and (x - 5/3)

OpenStudy (imstuck):

Got that?

OpenStudy (gazela1997):

yep and one more thing

OpenStudy (gazela1997):

did I factor this right? y=-3x-6x+9 y=-3(x+2x-3)

OpenStudy (anonymous):

\(\Large \color{blue}{3x^2+x-10=(x+2)(3x-5)}\)